OpenStudy (jkbo):

Please help me Which of the following equations is parallel to the given line? : y = -5x - 8 A.5x + y = 8 B.5x - y = 8 C.x + 5y = -8 D.-5x + y = -8

OpenStudy (anonymous):

if it is parallel then it has the same slope.

OpenStudy (anonymous):

however the answers are corrupted: x and y are on the same side!

OpenStudy (anonymous):

5x + y = 8 do you know how to fix this function ?

OpenStudy (jkbo):

@phanta_seea no I dont know how to

OpenStudy (anonymous):

you can move variables to the other sides if you change the sign: 5 x + y = +8 y = -5x +8

OpenStudy (anonymous):

the 5x was moved to the other side and changed sign. can you see?

OpenStudy (jkbo):

o ok so it A.

OpenStudy (anonymous):

yep :)

OpenStudy (jkbo):

thnak you

OpenStudy (jkbo):

@phanta_seea can you help me with one more? Which of the following equations is parallel to the given line? : 2x - 3y = 4  y = -2/3x + 4/3 y = -2x + 3 y = 3x - 2 y = 2/3x + 5

OpenStudy (anonymous):

ok, this time the line they have given is "corrupted"

OpenStudy (anonymous):

2 steps are necessary to get a proper function. the first one is similar to before - can you guess? 2x - 3y = 4

OpenStudy (anonymous):

ok watch and enjoy!! 1) y and x are on the same side, this shouldnt be! 2x -3y = 4 ....move -3y because we want y positive: 2x = 4 + 3y (sign of 3y has changed) 2) move 4 over to the other side: 2x = 4 + 3y 2x - 4 = 3y (notice how 4 sign has changed when switching sides!). 3) divide every term by 3: 2/3 x - 4/3 = y (note: divide every term! x term, coefficient and y)

OpenStudy (jkbo):

ok I got D. which is the last answer @phanta_seea

OpenStudy (anonymous):

its D

OpenStudy (jkbo):

Thank you again @phanta_seea